There is a 70 W heat leak from a room at temperature 22 °C into

an ideal refrigerator. How much electrical power is needed to

keep the refrigerator at -10 °C?


1
Expert's answer
2022-04-14T15:45:54-0400

"\frac{T_2}{T_1}=\frac{P_2}{P_1},"

"P_1=P_2\frac{T_1}{T_2}=78.5~W."